OpenStudy (jamanorthi):

how do i write 7.6% as a fraction in simplest form and as a decimal?

OpenStudy (anonymous):

7.6/100=76/1000=0.076

OpenStudy (jamanorthi):

how u get 76/1000

OpenStudy (anonymous):

7.6=76/10

OpenStudy (jamanorthi):

then

OpenStudy (anonymous):

76/10*100=76/1000

OpenStudy (jamanorthi):

thanks for helping

OpenStudy (anonymous):

:)

OpenStudy (jamanorthi):

how do i write 3.148% as a fraction in simplest form and as a decimal?

OpenStudy (anonymous):

you try it the same manner

OpenStudy (anonymous):

if percentage is given for any term like a% = a/100

OpenStudy (anonymous):

For your original questions this might help: Think of it this way 100=1.000 70=0.700 7=0.070